Twysted:

Need help with 4 questions, images in replies

Twysted:

1 attachment
Twysted:

1 attachment
Twysted:

1 attachment
Twysted:

1 attachment
rthrth:

i would use pythagorean theorem for a most of these \[a^2+b^2=c^2\] where a and b are two sides, c is the hypotenuse

rthrth:

dm me if you still need help

Manny300303199:

Crazy

umm:

First one: Setting up a proportion: \( \frac{77}{87.5} = \frac{22}{y} \) Solving for y, you found \( y = 25 \). Then, to find x, you added y to the given 22: \( x = 22 + 25 = 47 \) feet.

Manny300303199:

@umm wrote:
First one: Setting up a proportion: \( \frac{77}{87.5} = \frac{22}{y} \) Solving for y, you found \( y = 25 \). Then, to find x, you added y to the given 22: \( x = 22 + 25 = 47 \) feet.
You are suppose to guide the user not give them the answer. jp don't kill me jim

surjithayer:

for second triangles MAB and MNP are similar. \[\frac{ MB }{ MP }=\frac{ MA }{ MN }\] \[\frac{ x }{ 72.6 }=\frac{ 46.2-14 }{ 46.2 }\] \[\frac{ x }{ 72.6 }=\frac{ 32.2 }{ 46.2 }\] find x